Madge is six years older than Bruce. In five years the sum of their ages will be thirty-six. How is bruce
Found 2 solutions by ikleyn, TeachMath:
Answer by ikleyn(52787)

.
M - B = 6            ("Madge is six years older than Bruce")

(M+5) + (B+5) = 36   ("In five years . . . ")   =======>  M + B = 36 - 5 - 5 = 26.


So, you have 

M - B =  6
M + B = 26
--------------------  Add the equations

2M    = 6+26 = 32  ====>  M = 32/2 = 16.

==========>  B = 10.


ANSWER.  M is 16,  B is 10.



Answer by TeachMath(96)

If in 5 years' time, their ages sum to 36, then currently their ages sum to 36 - 2(5), or 36 - 10 = 26

Let Bruce's age be B
Then Madge's is: B + 6
We then get: B + B + 6 = 26
2B = 20
B, or Bruce is 20/2, or 10 years-old